OpenStudy (anonymous):

y = -x + 3 y = x - 3 A. (0, 3) B. (-3, 0) C. (3, 0) D. (0, -3) im a bit confused with it being a negative. can someone just walk me through with A and then with the negatives so i get it?

OpenStudy (anonymous):

What are your directions? Is it elimination?

OpenStudy (anonymous):

What is the solution to the system of equations graphed below? it basically has the lines and the equation thingy.. But that's all of it.. @JuanitaM

OpenStudy (anonymous):

You can find solution either via elimination or substitution. In this case eliminate.

OpenStudy (anonymous):

you can eliminate A

OpenStudy (displayerror):

You have the equations for two lines. See how both are equal to y? That means you can set them equal to each other: \[y = -x + 3 = x - 3\] We can then solve for x: \[-x + 3 = x-3\] Move the x terms to one side and the numbers to the other side: \[2x = 6\] Now solve for x. Once you have x, plug into either of these equations to solve for y: \[y = -x + 3 = x - 3\]

OpenStudy (anonymous):

(3,0)

OpenStudy (displayerror):

Yep!